# The influence of Parker spiral on the reflection-driven turbulence

### v1

@@ 5.2 Switchbacks
Magnetic switchbacks (SBs) are sharp, transient reversals of the heliospheric magnetic field direction that occur with little change in field magnitude. First detected by Helios and later observed in abundance by PSP, they are now recognized as a ubiquitous feature of the near-Sun solar wind (Bale2019; Kasper2019; Raouafi2023). SBs are typically highly Alfvénic, their magnetic and velocity fields remain strongly correlated, and $|\bm{B}|$ is nearly constant throughout the reversal (Bale2019).
Figure 12 illustrates representative fly-throughs at different stages of the evolution for both the radial and PS runs. In both geometries, $|\bm{B}|$ remains nearly constant while the field components undergo sharp directional swings in $B_{\parallel}$, accompanied by correlated deflections in $B_{T}$ and $B_{N}$. These signatures are consistent with large-amplitude, spherically polarized Alfvénic fluctuations as described in the past studies with similar setups (Squire2020; Shoda2021; Johnston2022). When $\delta B_{\parallel}$ reaches amplitudes comparable to the background, the local field direction can flip; the near constancy of $|\bm{B}|$ couples these radial reversals to transverse deflections, so each reversal appears as a coordinated multi-component swing. These rotations become noticeably sharper with expansion—compare the steeper swings between Fig. 12a and Fig. 12c—a natural consequence of the growing fluctuation amplitude. Furthermore, rotations in the PS runs are generally sharper than in the radial runs, as predicted by Squire2022. At late times, the radial case exhibits the loss of coordinated, constant-$|\bm{B}|$ rotations and increasingly irregular field-strength variations. These variations grow with distance, indicating that the magnetic field becomes less Alfvénic and more structurally complex as the cascade weakens. This late-time change broadly consistent with recent observational reports of enhanced field-strength variability with increasing heliocentric distance (Horbury2023). In contrast, the PS configuration retains a higher degree of Alfvénicity at comparable expansion factors, maintaining near-constant $|\bm{B}|$ through large directional rotations and better preserving its nearly spherically polarized character.
@@
Here, $z=0$ corresponds to a field perfectly aligned with the background, and $z=1$ to antiparallel configuration. We analyze regions corresponding to deflection thresholds $z_{\rm th}=0.125,\;0.25,\;0.375,\;0.5,\;0.625,\;0.75$. We find that PS runs produce a larger fraction of strong directional reversals ($z\gtrsim 0.5$), but a smaller fraction of weak deflections, than the corresponding radial runs (Fig. 13), in agreement with the 3D simulations of Johnston2022 and analytic explanation of Squire2022. In the PS case, the switchback fraction grows until a near spiral angle of $45^{\circ}$ and then declines. This decline coincides with the change in the evolution of $v_{\rm A}$ between the two geometries discussed in section˜2. The right panel of Fig. 13 shows the evolution of the normalized amplitude $z^{+}/v_{\rm A}$ for all the simulation sets which are compared to the WKB expectations for a radial ($z^{+}/v_{\rm A}\propto a^{1/2}$) or azimuthal ($z^{+}/v_{\rm A}\propto a^{-1/2}$) background field. In the radial case, the amplitude rise slowly with increasing $a$, whereas the PS runs show a different behavior: they remain approximately flat or grow slowly at small $a$ and then decline at large $a$.
@@ 5.3 Compressibility
$C_{B}$ is nearly zero for pure transverse waves and grows when compressive fluctuations are present (Chen2021; Shoda2021). $\delta\rho_{\rm rms}/\langle\rho\rangle$ measures the fractional amplitude of density variations relative to the mean density. Taken together, $C_{B}$ and $\delta\rho_{\rm rms}$ measure the prevalence of compressive fluctuations in the system. Figure 14 shows the evolution of $C_{B}$, and $\delta\rho_{\rm rms}/\langle\rho\rangle$ as functions of the expansion factor $a$. All runs start with higher $C_{B}$ because of the initialization of simulations with large-amplitude linearly polarized waves. As expansion and nonlinear mode coupling proceed, compressions in $|\bm{B}|$ rapidly dissipated within $\leq 1\tau_{\rm A}$, driving the system toward spherical polarization and an effectively constant $|\bm{B}|$; correspondingly, $C_{B}$ drops significantly at early times.
@@
We note an important caveat: the simulations presented here employ an isothermal equation of state, which does not capture the solar wind’s full thermodynamics. A more realistic equation of state should be included for direct quantitative comparison with spacecraft density signatures and compressive energetics.
@@ 6 Conclusion
With a purely radial mean field the simulations show two phases. At small heliocentric distances the flow is strongly imbalanced $(z^{+}\gg z^{-})$ and highly Alfvénic, while at larger radii the inward component grows, and the system approaches a more balanced state $(z^{+}\sim z^{-})$. The dominant $z^{+}$ component undergoes both forward (to smaller scales) and inverse (to larger scales) transfer, and the net result is a weakening of nonlinear interactions as $\chi_{\rm exp}$ approaches 1. The phenomenology developed in section˜2.4, which is effectively standard one of Dmitruk2002, predicts $z^{+}/z^{-}\sim\chi_{\rm exp}$; hence $\chi_{\rm exp}\sim 1$ marks the end of the imbalanced phase. During this late phase the Elsässer fields become strongly aligned $(\bm{z}^{-}\propto-\bm{z}^{+})$, the flow organizes into quasi two-dimensional, magnetically dominated Alfvénic vortices, and volumetric dissipation is reduced. Our results are broadly consistent with the reduced MHD study of Meyrand2025, extending them to a compressible framework. This also shows how constant-$|\bm{B}|$ spherically polarized switchbacks give way to more compressive balanced turbulence with larger $\lvert\bm{B}\rvert$ variation, a transition also observed in the solar wind (Horbury2023).
@@
We provide various additional diagnostics that could be interesting to connect to in-situ observations. The joint evolution on the circle plot $(\sigma_{c},\sigma_{r}$; Fig. 11)—from high $\sigma_{c}$, small $\sigma_{r}$ near the Sun toward reduced $\sigma_{c}$ and more negative $\sigma_{r}$ at larger radii—depends on the Parker spiral angle (Bavassano1998; Bruno2007; Wicks2013; Meyrand2025), with the PS making it stray further from the circles edge. Interestingly, the PS geometry does not systematically increase or decrease compressive activity; compressibility diagnostics (e.g., $C_{B}$, $\delta\rho_{\rm rms}/\langle\rho\rangle$) show no persistent enhancement in PS runs. Rather, the differences between PS and radial cases reflect that PS intervals remain more Alfvénic and, as a consequence, preserve near constant-$|\bm{B}|$ rotations to larger $a$. This yields generally sharper, more Alfvénic switchbacks for the PS geometry than in a purely radial field. Including a Parker spiral component tends to increase the growth and persistence of switchbacks with expansion, but the slower decay of $v_{\rm A}$ once $\Phi>45^{\circ}$ can halt that growth (Fig. 13). These predictions could be directly testable with spacecraft observations, perhaps via comparisons across heliocentric latitudes.
@@
Our results are subject to several controlled approximations: they are obtained from compressible isothermal MHD simulations in an expanding-box framework with constant solar-wind speed. The model therefore applies most directly outside the Alfvén point for the parameter ranges explored here—processes that amplify fluctuations inside the Alfvén point (and the accelerating background) are not included and there is no PS in such a regime anyway. Accordingly, our simulations should be interpreted as the nonlinear evolution of large-amplitude outward $z^{+}$ perturbations that are already present at or beyond the Alfvén point. To study the early-stage amplification of waves, one must adopt frameworks that capture the accelerating wind (Chandran2019; Ballegooijen2011; Tenerani2017; Shoda2021).
@@
In addition to these caveats, we note several important limitations and directions for future work. First, the finite box size in the parallel direction ($L_{\parallel}$) constrains the longest-parallel wavelengths that can be represented and can therefore bias measured correlation lengths and decay rates; systematic domain-size scans are needed to quantify these effects. Second, the isothermal MHD approximation omits kinetic physics (e.g., collisionless damping, pressure anisotropy, and other wave-particle processes) that can modify both reflection and dissipation, so extensions to kinetic or hybrid models are desirable to assess how robust the present MHD results are. Third, some aspects of our conclusions depend on the assumed spectral parameters of the initial conditions (e.g. $k_{\rm width}$ and $\kappa$); the sensitivity of the dynamics to the initial spectral width should be explored with controlled variations. Finally, we have not fully explored the space of expansion-nonlinearity ratios: further runs with different $\chi_{\rm exp}$ and with higher initial fluctuation amplitudes are required to map the transition to stronger nonlinearity and to determine how the results scale with initial conditions.

### v2

@@ 5.2 Switchbacks
Magnetic switchbacks (SBs) are sharp, transient reversals of the heliospheric magnetic field direction that occur with little change in field magnitude. First detected by Helios and later observed in abundance by PSP, they are now recognized as a ubiquitous feature of the near-Sun solar wind (Bale et al., 2019; Kasper et al., 2019; Raouafi et al., 2023). SBs are typically highly Alfvénic, their magnetic and velocity fields remain strongly correlated, and $|\bm{B}|$ is nearly constant throughout the reversal (Bale et al., 2019).
Figure 12 illustrates representative fly-throughs at different stages of the evolution for both the radial and PS runs. In both geometries, $|\bm{B}|$ remains nearly constant while the field components undergo sharp directional swings in $B_{\parallel}$, accompanied by correlated deflections in $B_{T}$ and $B_{N}$. These signatures are consistent with large-amplitude, spherically polarized Alfvénic fluctuations as described in the past studies with similar setups (Squire et al., 2020; Shoda et al., 2021; Johnston et al., 2022). The emergence of nearly constant-$|\bm{B}|$ fluctuations reflects the nonlinear evolution’s tendency to favor spherically polarized, Alfvénic states that preserve field magnitude (Barnes and Hollweg, 1974). This is likely because any perturbation that changes $|\bm{B}|$ excites compressive (fast or slow magnetosonic) modes with associated density and pressure variations; these compressive components steepen into shocks and dissipate more rapidly than the incompressible Alfvénic part that maintains constant field magnitude.
When $\delta B_{\parallel}$ reaches amplitudes comparable to the background, the local field direction can flip; the near constancy of $|\bm{B}|$ couples these radial reversals to transverse deflections, so each reversal appears as a coordinated multi-component swing. These rotations become noticeably sharper with expansion—compare the steeper swings between Fig. 12a and Fig. 12c—a natural consequence of the growing fluctuation amplitude. Furthermore, rotations in the PS runs are generally sharper than in the radial runs, as predicted by Squire et al. (2022). At late times, the radial case exhibits the loss of coordinated, constant-$|\bm{B}|$ rotations and increasingly irregular field-strength variations. These variations grow with distance, indicating that the magnetic field becomes less Alfvénic and more structurally complex as the cascade weakens. This late-time change broadly consistent with recent observational reports of enhanced field-strength variability with increasing heliocentric distance (Horbury et al., 2023). In contrast, the PS configuration retains a higher degree of Alfvénicity at comparable expansion factors, maintaining near-constant $|\bm{B}|$ through large directional rotations and better preserving its nearly spherically polarized character.
@@
Here, $z=0$ corresponds to a field perfectly aligned with the background, and $z=1$ to antiparallel configuration. We analyze regions corresponding to deflection thresholds $z_{\rm th}=0.125,\;0.25,\;0.375,\;0.5,\;0.625,\;0.75$. We find that PS runs produce a larger fraction of strong directional reversals ($z\gtrsim 0.5$), but a smaller fraction of weak deflections, than the corresponding radial runs (Fig. 13), in agreement with the 3D simulations of Johnston et al. (2022) and analytic explanation of Squire et al. (2022). In the PS case, the switchback fraction grows until a near spiral angle of $45^{\circ}$ and then declines. This decline coincides with the change in the evolution of $v_{\rm A}$ between the two geometries discussed in section˜2. The right panel of Fig. 13 shows the evolution of the normalized amplitude $z^{+}/v_{\rm A}$ for all the simulation sets which are compared to the WKB expectations for a radial ($z^{+}/v_{\rm A}\propto a^{1/2}$) or azimuthal ($z^{+}/v_{\rm A}\propto a^{-1/2}$) background field. In the radial case, the amplitude rise slowly with increasing $a$, whereas the PS runs show a different behavior: they remain approximately flat or grow slowly at small $a$ and then decline at large $a$.
@@ 5.3 Compressibility
$C_{B}$ is nearly zero for pure transverse waves and grows when compressive fluctuations are present (Chen et al., 2021; Shoda et al., 2021). $\delta\rho_{\rm rms}/\langle\rho\rangle$ measures the fractional amplitude of density variations relative to the mean density. Taken together, $C_{B}$ and $\delta\rho_{\rm rms}$ measure the prevalence of compressive fluctuations in the system. Figure 14 shows the evolution of $C_{B}$, and $\delta\rho_{\rm rms}/\langle\rho\rangle$ as functions of the expansion factor $a$. All runs start with higher $C_{B}$ because of the initialization of simulations with large-amplitude linearly polarized waves. As expansion and nonlinear mode coupling proceed, compressions in $|\bm{B}|$ rapidly dissipated within $\leq 1\tau_{\rm A}$, driving the system toward spherical polarization and an effectively constant $|\bm{B}|$; correspondingly, $C_{B}$ drops significantly at early times.
@@
We note an important caveat: the simulations presented here employ an locally isothermal equation of state, which does not capture the solar wind’s full thermodynamics. A more realistic equation of state should be included for direct quantitative comparison with spacecraft density signatures and compressive energetics.
@@ 6 Conclusion
With a purely radial mean field the simulations show two phases. At small heliocentric distances the flow is strongly imbalanced $(z^{+}\gg z^{-})$ and highly Alfvénic, while at larger radii the inward component grows, and the system approaches a more balanced state $(z^{+}\sim z^{-})$. The dominant $z^{+}$ component undergoes both forward (to smaller scales) and inverse (to larger scales) transfer, and the net result is a weakening of nonlinear interactions as $\chi_{\rm exp}$ approaches 1. The phenomenology developed in section˜2.4, which is effectively standard one of Dmitruk et al. (2002), predicts $z^{+}/z^{-}\sim\chi_{\rm exp}$; hence $\chi_{\rm exp}\sim 1$ marks the end of the imbalanced phase. During this late phase the Elsässer fields become strongly aligned $(\bm{z}^{-}\propto-\bm{z}^{+})$, the flow organizes into quasi two-dimensional, magnetically dominated Alfvénic vortices, and volumetric dissipation is reduced. Our results are broadly consistent with the reduced MHD study of Meyrand et al. (2025), extending them to a compressible framework. This also shows how constant-$|\bm{B}|$ spherically polarized switchbacks give way to more compressive balanced turbulence with larger $\lvert\bm{B}\rvert$ variation, a transition also observed in the solar wind (Horbury et al., 2023).
@@
We provide various additional diagnostics that could be interesting to connect to in-situ observations. The joint evolution on the circle plot $(\sigma_{c},\sigma_{r}$; Fig. 11)—from high $\sigma_{c}$, small $\sigma_{r}$ near the Sun toward reduced $\sigma_{c}$ and more negative $\sigma_{r}$ at larger radii—depends on the Parker spiral angle (Bavassano et al., 1998; Bruno et al., 2007; Wicks et al., 2013; Meyrand et al., 2025), with the PS making it stray further from the circles edge. Interestingly, the PS geometry does not systematically increase or decrease compressive activity; compressibility diagnostics (e.g., $C_{B}$, $\delta\rho_{\rm rms}/\langle\rho\rangle$) show no persistent enhancement in PS runs. Rather, the differences between PS and radial cases reflect that PS intervals remain more Alfvénic and, as a consequence, preserve near constant-$|\bm{B}|$ rotations to larger $a$. This yields generally sharper, more Alfvénic switchbacks for the PS geometry than in a purely radial field. Including a Parker spiral component tends to increase the growth and persistence of switchbacks with expansion, but the slower decay of $v_{\rm A}$ once $\Phi>45^{\circ}$ can halt that growth (Fig. 13). These predictions could be directly testable with spacecraft observations, perhaps via comparisons across heliocentric latitudes.
@@
Our results are subject to several controlled approximations: they are obtained from compressible locally isothermal MHD simulations in an expanding-box framework with constant solar-wind speed. The model therefore applies most directly outside the Alfvén point for the parameter ranges explored here—processes that amplify fluctuations inside the Alfvén point (and the accelerating background) are not included and there is no PS in such a regime anyway. Accordingly, our simulations should be interpreted as the nonlinear evolution of large-amplitude outward $z^{+}$ perturbations that are already present at or beyond the Alfvén point. To study the early-stage amplification of waves, one must adopt frameworks that capture the accelerating wind (Chandran and Perez, 2019; van Ballegooijen et al., 2011; Tenerani and Velli, 2017; Shoda et al., 2021).
@@
In addition to these caveats, we note several important limitations and directions for future work. First, the finite box size in the parallel direction ($L_{\parallel}$) constrains the longest-parallel wavelengths that can be represented and can therefore bias measured correlation lengths and decay rates; systematic domain-size scans are needed to quantify these effects. Second, the locally isothermal MHD approximation omits kinetic physics (e.g., collisionless damping, pressure anisotropy, and other wave-particle processes) that can modify both reflection and dissipation, so extensions to kinetic or hybrid models are desirable to assess how robust the present MHD results are. Third, some aspects of our conclusions depend on the assumed spectral parameters of the initial conditions (e.g. $k_{\rm width}$ and $\kappa$); the sensitivity of the dynamics to the initial spectral width should be explored with controlled variations. Finally, we have not fully explored the space of expansion-nonlinearity ratios: further runs with different $\chi_{\rm exp}$ and with higher initial fluctuation amplitudes are required to map the transition to stronger nonlinearity and to determine how the results scale with initial conditions.
Acknowledgements
@@
Declaration of interest
The authors report no conflict of interest.